Credit risk impacts WACC: True / False
Answer #1

True.

Credit risk impacts WACC.

As Cost of debt depends on Credit Risk, so it impact WACC.

First of all, WACC stands for Weighted Average Cost of Capital and it calculated using following formula.

WACC = We*Ke + Wd*Kd*(1-Income Tax rate)

We = Weight of Equity = E/E+D (Equity portion of total financing)

(Where E is market value total equity)

Re = total cost of equity

Wd = D/E+D (debt portion of total financing)

(Where D is market value of total debt)

Rd = total cost of debt

Cost of Debt is nothing but interest paid on funds , when companies borrow funds from lenders.

Basically the Cost of Debt = Kd = Rf+Credit Risk rate

(The cost of debt is computed by taking the rate on a risk-free bond whose duration matches the term structure of the corporate debt, then adding a default premium)

Default premium is nothing but Credit risk rate.

Hence its true that Credit risk impacts WACC as it impact Kd which is nothing but Cost of Debt.